OpenStudy (omgthatgirl):

Find the center and radius of the circle with equation (x-6)^2 + (y+5)^2 =9 A. Center (6,-5); radius 3 B. Center (-6,5); radius 9 C. Center (6,-5); radius 9 D. Center (-6,5); radius 3

rishavraj (rishavraj):

EQUATION OF CIRCLE WITH CENTER (h , k) AND RADIUS r \[(x - h)^2 + (y - k)^2 = r^2\]

OpenStudy (omgthatgirl):

ok will the radius be 9 @rishavraj

OpenStudy (freckles):

r^2=9 r is not also 9

OpenStudy (freckles):

if r^2=9 then r=?

OpenStudy (omgthatgirl):

3

OpenStudy (freckles):

yep

OpenStudy (omgthatgirl):

so the answer is a

OpenStudy (freckles):

sounds great

OpenStudy (omgthatgirl):

ok thank you